ecplain why x and x+1 cannot represent two consecutive odd integers
Answer by jim_thompson5910(35256)

Let's say x is odd. That would make x+1 even.

example:

x = 3

x+1 = 3+1 = 4

so we see that if x = 3, then x+1 = 4. The first number is odd, but the next number is even.

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------